OpenStudy (asheffield):

f(x) = 1 + 3bx + 2x^2 x =< 1 f(x) = mx+b x > 1 Find the values of m and b so that f will be both continuous and differentiable at x=1 I got m = 1 and b = -1, but my work wasn't pretty and I'm not sure I believe my answer.

OpenStudy (hafeda):

For f to be continuous at x = 1, we must have the two parts of the function equal, so 1 + 3bx + 2x^2 = mx + b at x = 1: 1 + 3b + 2 = m + b 3b + 3 = m + b 2b = m - 3. first equation [1] For the function to be differentiable, it means the derivatives of the two parts must match. So: f '(x) = {3b + 4x or m} So we have 3b + 4x = m, at x = 1; 3b + 4 = m. second equation [2]. Solve [1] and [2] by substitution or solve for b or m and sub then: [1] 2b = m - 3 [2] m = 3b + 4 so 2b = (3b + 4) - 3 2b = 3b + 1 -b = 1 b = -1, and with [2] m = 3b + 4 = 3*(-1) + 4 = 1. So the answer is (m, b) = (1, -1)